the area of a rectangle is 300 square inches. the sides are in a 3:4 ratio. what is the perimeter of the rectangle?
Answer by richwmiller(17219)

3x*4x=300
12x^2=300
x^2=25
x=5
p=2*(3x+4x)=14x=70